Session: Feb 2012 (session 1)
Wavelength: 1.3 cm


To give feedback on this experiment, click here

Experiment Feedback

Yebes (Angel Vicente at Tue Feb 28 22:36:23 2012) Success:
No problem

Hartebeesthoek (Jonathan Quick at Tue Feb 28 22:40:54 2012) Minor failures: FS
Some IFD communication problems. FS (antcn) hang caused data loss from 19:28-19:34 UT

Urumqi (Wenjun Yang at Tue Feb 28 23:01:38 2012) Success:

Shanghai (Bo Xia at Tue Feb 28 23:12:13 2012) Success:

Svetloe (Slobojaninova N. at Wed Feb 29 11:27:39 2012) Success:
Scan No0132 - failed due to antenna problems

Zelenchukskaya (Kononenko V. at Wed Feb 29 11:28:21 2012) Success:
Scans: No0022 - No0061 failed due antenna problems

Effelsberg (Uwe Bach at Thu Mar 01 09:26:06 2012) Minor failures:RECE
Lost IF power from 17-19 UT. Reason unknown, maybe some misbehaving switch, restart of the frontend solved the problem.

Onsala (Michael Lindqvist at Mon Mar 05 10:55:04 2012) Success:
No known problems.

Metsahovi (E. Rastorgueva at Mon Mar 05 11:49:57 2012) Minor failures: WIND
Very strong snow storm during the observation, layer of snow on the radome. By the end of an experiment started to melt the snow.

Jodrell_Bank (Alastair Gunn at Mon Mar 19 15:13:14 2012) Success:
No known problems.